OpenStudy (anonymous):

mardi received an inheritance of $60.000. she invested part at 10% and deposited the remainder in tax-free bonds at 12%. her total annual income from the investment was $6400. find the amount invested at 10%

OpenStudy (baru):

x=money invested for 10% y= money invested for 12% x+y=60000 (x*10/100)+(y*12/100)=6400

OpenStudy (anonymous):

thank you!